### v2

# Layer-Wise Quantization: A Pragmatic and Effective Method for Quantizing LLMs Beyond Integer Bit-Levels

Abstract
We present a simple variable quantization approach that quantizes different layers of a large language model (LLM) at different bit levels. Specifically, we quantize the most important layers to higher bit precision and less important layers to lower bits to achieve floating point quantization levels. We propose two effective strategies to measure the importance of layers within LLMs: the first measures the importance of a layer based on how different its output embeddings are from the input embeddings (the higher the better); the second estimates the importance of a layer using the number of layer weights that are much larger than average (the smaller the better). We show that quantizing different layers at varying bits according to our importance scores results in minimal performance drop with a far more compressed model size. Finally, we present several practical key takeaways from our variable layer-wise quantization experiments: (a) LLM performance under variable quantization remains close to the original model until 25-50% of layers are moved in lower quantization using our proposed ordering but only until 5-10% if moved using no specific ordering; (b) Quantizing LLMs to lower bits performs substantially better than pruning unless extreme quantization (2-bit) is used; and (c) Layer-wise quantization to lower bits works better in the case of larger LLMs with more layers compared to smaller LLMs with fewer layers. The code used to run the experiments is available at: https://github.com/RazvanDu/LayerwiseQuant.
11footnotetext: The work was done while Razvan-Gabriel Dumitru interned at ServiceNow.
Layer-Wise Quantization: A Pragmatic and Effective Method for Quantizing LLMs Beyond Integer Bit-Levels
Razvan-Gabriel Dumitru1,2, Vikas Yadav1, Rishabh Maheshwary1, Paul-Ioan Clotan3, Sathwik Tejaswi Madhusudhan1, Mihai Surdeanu2 1ServiceNow Research, 2University of Arizona, 3Università di Bologna Correspondence: razvandumm@gmail.com
1 Introduction
Figure 1: Overall intuition behind our approach. We first rank the layers in an LLM (e.g., LLaMa-2-13B) using an importance score (shown here is ranking based on our Layer Input Modification (LIM) score, see section 3.1). The color intensity in each layer represent their LIM importance score. Then, the 30 most important layers are quantized in 4 bit while the remaining 10 least important layers are quantized in 2 bits, resulting in 3.5 bits as the average bit size.
Large Language Models (LLMs) have achieved remarkable performance on a variety of tasks, especially when scaled to billions of parameters Jiang et al. (2023, 2024); Touvron et al. (2023); Zhang et al. (2022); Team et al. (2023). The largest open-source models available can have an upwards of 400B parameters, such as LLaMa3-400B Touvron et al. (2023). Even small models such as LLaMa3-8B require as much as 20GB of VRAM to run on a GPU at the original precision, making them unusable for low resource settings. In such settings model compression techniques such as quantization are critical Zhu et al. (2023b); Wan et al. (2023).
Most prominent techniques for model compression broadly cover pruning Ma et al. (2023), knowledge distillation Gu et al. (2023), and quantization Zhu et al. (2023b). Pruning yields improvements in inference speed, but often results in substantial performance drop Frantar and Alistarh (2023); Men et al. (2024). On the other hand, quantization has proven to be a more robust solution for model size compression with comparatively much smaller performance drops Lin et al. (2024). In our work, we primarily focus on memory reduction through quantization. Further, quantization can be training specific or post-training Yao et al. (2024), where trained models are quantized without requiring any further training. We focus on post-training quantization due to its practicality.
The majority of quantization techniques proposed recently (Frantar et al., 2023; Xiao et al., 2023; Yao et al., 2022, inter alia) focus on the quantization of all the LLM layers to a single precision bit, or quantizing only specific weights of the network Lin et al. (2024). This has shown to be costly, their effectiveness is data dependent, and their implementations can be considered moderately challenging. In contrast, we propose a simple technique that quantizes different layers at different bit precision depending on their importance. Figure 1 shows an example of our idea. We show that our approach is simple to implement, performs well, and provides greater flexibility to compress models in varying bit precision as per memory requirements. Further, our approach is complementary to other quantization techniques. For example, in our experiments we coupled our idea with two effective and prominent quantization techniques: GPT-Q and Quanto, but any other techinques can be used.
The contributions of our work are as follows:
Figure 2: Plots showing the effect of variable quantization for LLaMa2-13b and multiple datasets using Quanto. The leftmost point indicates LLM performance when all 40 layers of the LLM are represented in 4-bit; the rightmost point shows LLM performance when all layers are quantized to 2-bit. The dots on each curve (in each plot) show accuracy when the model is quantized to lower bits by converting less important layers to 2 bits one by one. Red and purple line indicate performance from 8bit and fp16 precision model (ceiling models). As shown, there is no considerable performance drop from fp16 or 8bit to 4bit precision. Hence, we focus our experiments on quantization below 4 bits. The vertical gray line indicates the quantization point that preserves 90% of the 4bit performance. The red line represents when layers are ordered randomly. We chose 3 random orders of the layers and quantized layers to 2 bits as per these orders. The standard deviation in performance from random orders are highlighted on the red curve. The curves are plotted on 2K evaluation data while results on full data is summarized in table 1.
Figure 3: Plots showing the effect of variable quantization for Mistral-7b and multiple datasets using Quanto. All notations are the same as in Figure 2.
  • We introduce a novel method for layer-wise quantization of LLMs at different bit levels to achieve flexible lower bit precision overall, enabling LLMs to fit into range of specified memory capacity. We accompany our method with a detailed study, which highlights various important findings such as importance of layer ranking for quantizing less important layers with lower bits and more important layers with higher bit precision. We show that models can be quantized significantly more from 4-bit while retaining 90% of performance when following our proposed order.
  • We propose and study two layer importance scores for variable quantization. Our first score named layer input modification (LIM) is based on how much a layer changes its input representations into the output ones. This score is calculated with an unlabeled text calibration corpus. The second scoring method, called z-score distribution (ZD), measures the distribution of weight magnitudes within a layer to determine its importance. To our knowledge, we are the first to propose layer orderings for quantization. We validate these scores by empirically showing that when LLM layers are quantized to lower bits as per rankings from our two importance score, they retain performance much more strongly than several other layer ordering baselines.
  • We evaluate the impact of our variable quantization method based on layer importance on five top performing LLMs from different model families and different sizes. We draw several important practical findings from these experiments: (a) LLM performance under variable quantization remains close to the original model when using our ordering until reaching the level of 3.0-3.25 bits on average; (b) Quantizing LLMs to lower bits performs substantially better than pruning; however, under extreme quantization settings (i.e., $<=2-bits$) pruning shows slightly better results; (c) We show that quantization at two levels (i.e., quantizing some layers in $x$ bits and remaining layers in $y$ bits) performs much better than three levels of quantization, suggesting that the interaction between layers with different quantization is complex and may need more investigation; and (d) Layer-wise quantization to lower bits works better in the case of larger LLMs with more layers compared to smaller LLMs with fewer layers.
2 Related Work
Quantization techniques for large language models (LLMs) aim to reduce the precision of weights and activations to lower-bits without significantly compromising performance Zhu et al. (2023b). Popular approaches include Post-Training Quantization (PTQ) Banner et al. (2019) and Quantization-Aware Training (QAT) Liu et al. (2023). PTQ can be divided into static quantization, which uses a small dataset to calibrate scaling factors for weights and activations, and dynamic quantization, which quantizes activations on-the-fly during inference. Our study utilizes the static PTQ techniques such as GPT-Q in our experiments for practicality.
A few works have highlighted 4 bit precision as a robust quantization limit for wide variety of NLP tasks Dettmers and Zettlemoyer (2023). Our results (fig. 2 and 3) also show similar findings that performance drop is very minute from bf16 to 8 bits, and then to 4 bits. Hence, our experiments focus more on quantizing LLMs below 4 bits to highlight effectiveness of layer-wise quantization based on layer importance. Specifically, our empirical results (section 5) show that variable layer-wise quantization can retain 90% of the performance with a notable compression upto 2.85 bits overall. Concurrent (ArXiv preprint) work by Tai et al. (2024) also show effectiveness of variable quantization in vision language models.
A few previous works have studied layer importance in transformers Vaswani et al. (2017); Simoulin and Crabbé (2021). Some of the recent, (unpublished) concurrent works such as ShortGPT Men et al. (2024) have also proposed utilizing layer importance but primarily focusing on pruning. Shen et al. (2020) had utilized hessian information from each layer as an importance measure to quantize specific weight matrices at different bits. Different from these, our work is focused more on utilizing layer importance for quantizing more important layers (fully) in higher bits and less important layers in lower bits. Importantly, our proposed approach of layer-wise quantization of LLMs based on their layer importance is a method that can be easily extended with any quantization techniques like GPT-Q and Quanto (section 3.2). We also compare ShortGPT layer importance based pruning baseline with our layer-wise quantization approach in section 6.1.
3 Method
In our study, given a LLM with N layers denoted as $\{\mathbf{L_{1},L_{2},L_{3},....L_{N}}\}$, we first compute the importance of each layer, and then quantize them differently based on their importance. To achieve the former, we propose two scoring methods that allow us to rank the layers based on their respective importance scores.
3.1 Layer Importance Scores
We call our first layer importance score layer input modification (LIM). The intuition behind LIM is that the more a layer changes its received input embeddings the more important it must be. More formally, the LIM score for a specific layer $\mathbf{L_{i}}$ measures the modification of the representations of $\mathbf{L_{i}}$’s input (denoted by $\mathbf{L_{i}^{I}}$) to create its output representations (denoted by $\mathbf{L_{i}^{O}}$). A near similar score has been also studied for pruning (layer removal) in a concurrent (unpublished) work Men et al. (2024). The LIM score is the negative of cosine similarity between $\mathbf{L_{i}^{I}}$ and $\mathbf{L_{i}^{O}}$ as shown below:
$\vspace{-2mm}\text{LIM}(\mathbf{L_{i}})=-\frac{\mathbf{L_{i}^{I}}\cdot\mathbf{L_{i}^{O}}}{\|\mathbf{L_{i}^{I}}\|\|\mathbf{L_{i}^{O}}\|}\vspace{+1mm}$ (1)
The dot product $\mathbf{L_{i}^{I}}\cdot\mathbf{L_{i}^{O}}$ quantifies the alignment between the input and output vectors, while the normalization factor $\|\mathbf{L_{i}^{I}}\|\|\mathbf{L_{i}^{O}}\|$ scales the dot product to the range of [-1, 1], corresponding to the cosine of the angle between the two vectors. The negative sign in the LIM score indicates that a higher similarity (cosine similarity closer to 1) between $\mathbf{L_{i}^{I}}$ and $\mathbf{L_{i}^{O}}$ results in a lower importance score. For measuring change in $\mathbf{L_{i}^{I}}$ to $\mathbf{L_{i}^{O}}$, we use all 50 documents of pg19! Rae et al. (2019) (an unlabelled text corpus) as calibration corpus.
Our second score, called z-score distribution (ZD), is based on the distribution of parameter values and does not require any calibration data. Intuitively, this score considers a layer as more important if it contains more weights that are much higher than average.
We examine the proportion of weights in a layer exhibiting a z-score greater than 1. The z-score of a weight $w_{i}$ is defined as
$z_{i}=\frac{w_{i}-\mu}{\sigma},$
where for layer $\mathbf{L_{i}}$, $w_{i}$ represents an individual weight, $\mu$ the mean of the weights, and $\sigma$ their standard deviation. The measure of interest, $ZD(L_{i})$, is expressed as
$\text{ZD}(L_{i})=\frac{|Li_{Zscore}>1|}{N_{i}},$
quantifying the ratio of weights whose z-scores exceed 1 ($|Li_{Zscore}>1|$) to the total number of weights $N_{i}$ in the $i^{th}$ layer.
Layer ranking baselines:
To highlight the benefits of our layer importance scores, we compare them to two baselines. First, we randomly rank layers with different random seeds to highlight ordering of layer is important. Second, to specifically highlight the strengths of the LIM score, we consider a reverse LIM ordering as a baseline. Finally, we also considered the reverse index of layers, i.e., from layer 30 to layer 1, following Gromov et al. (2024) who has shown just removing layers from the top of the LLM to be an effective order for layer pruning. We implement the same pruning with specified reverse indexes from Gromov et al. (2024) and show its comparison in Figure 4(a).
3.2 Quantization Techniques
In our work, we employ two post-training quantization techniques—GPT-Q and Quanto—to our proposed layer-wise quantization study.
3.2.1 Quanto
Quanto is a fast-acting quantization method [^1] https://github.com/huggingface/optimum-quanto that simplifies the process of reducing precision after training. In practice, Quanto achieves quicker quantization by applying uniform scaling factors across all layers of a model, avoiding the need for detailed data-driven analysis of each layer’s distribution. Similarly to most techniques, it allows for int2, int4, int8, and fp8 quantization. In our implementation for variable quantization of different layers, we quantized the same model to two different bit levels (lets say int2 and int4). Then, based on the layer importance order, we selected less important layers from the int2 quantized layer sets and more important layers from int4 quantized sets.
[…]
Table 1: Accuracy on full evaluation datasets of different models quantized with Quanto. All layers start at 4 bits; we then quantize N number of layers in 2 bits where N is mentioned in the “Layers 2 bits” column. We also show results for 8 to 4 bit quantization in Appendix in table 4. Average performances within 90% of the 4 bit model are highlighted in bold.
GPT-Q Quantization
Model Avg. Layers WNGD ARC PIQA HLSWG MMLU Avg.Acc
bits 2 bits
LIM Ordering LLama-7B 3.68 5 67.1 73.4 76.2 54.9 37.5 61.8
3.37 10 67.6 67.5 73.4 52.9 35.2 59.3
3.06 15 65.1 56.0 68.0 46.4 31.5 53.4
Mistral-7B 3.68 5 73.1 77.4 78.0 59.1 55.4 68.6
3.37 10 70.6 74.8 76.9 56.8 54.9 66.8
3.06 15 66.1 65.3 72.9 50.7 42.0 59.4
LLama-13B 3.75 5 71.4 77.5 78.5 59.0 49.3 67.1
3.50 10 71.8 76.4 77.7 58.2 48.2 66.5
3.25 15 72.6 73.7 75.7 56.7 47.3 65.2
Qwen-2-7B 3.64 5 62.0 67.2 77.0 52.0 56.6 63.0
3.28 10 56.7 53.1 71.4 45.5 29.4 51.2
2.92 15 53.4 45.0 66.4 42.0 25.0 46.4
Z-score Ordering LLama-7B 3.68 5 67.2 71.1 75.9 54.9 38.1 61.4
3.37 10 68.0 66.9 73.1 52.1 33.7 58.8
3.06 15 62.9 56.7 67.8 46.2 28.5 52.4
Mistral-7B 3.68 5 72.9 77.7 78.8 59.2 55.1 68.7
3.37 10 69.1 72.1 74.8 53.2 38.9 61.6
3.06 15 62.0 52.7 65.6 39.5 25.1 49.0
LLama-13B 3.75 5 - - - - - -
3.50 10 71.3 75.0 76.4 57.4 48.2 65.6
3.25 15 72.3 73.5 76.2 56.4 45.7 64.8
Qwen-2-7B 3.64 5 69.1 71.1 75.2 54.0 64.3 66.8
3.28 10 65.6 65.2 71.6 48.2 47.5 59.6
2.92 15 54.8 50.0 66.7 42.8 30.4 49.4
Table 2: Comparison of different models and their performance across various tasks with GPT-Q quantization.
3.2.2 GPT-Q
GPT-Q is a post-training quantization technique specifically designed for GPT models Frantar et al. (2023). Utilizing a dataset, it calculates the necessary scaling factors for quantization. After training, GPT-Q assesses the distribution of weights using this dataset to determine optimal scaling factors for converting floating-point representations to lower-bit formats such as int8 or int4. We only have several data points for GPT-Q and it adds significant execution time overhead to our experiments. Both Quanto and GPT-Q are available in HuggingFace library Wolf et al. (2019). We modified it for our use-case in the same way as described for Quanto.
4 Experiments
4.1 Models and Hyperparameters
We studied quantization on 5 LLMs from different model families and different sizes – LLaMa-2-7b Touvron et al. (2023), LLaMa-2-13B, Mistral-7b Jiang et al. (2023), QWEN-1.8b, and QWEN-7b Bai et al. (2023). We evaluated these LLMs on 8 A100 GPUs with a batch size of 1 for inference using LLM harness library Gao et al. (2023).
4.2 Evaluation Datasets
We select five diverse NLP tasks for evaluating the quantization effects: Winogrande Sakaguchi et al. (2021), ARC-easy Clark et al. (2018), PIQA Bisk et al. (2020), HellaSwag Zellers et al. (2019), and MMLU Hendrycks et al. (2020). We also evaluate our approach on two generation datasets to cover diverse tasks of reasoning and answer generation: GSM8K Cobbe et al. (2021), which contains math questions, and the Natural Questions (open) dataset Kwiatkowski et al. (2019), which consists of open-domain answer generation task.
5 Discussion of Results
Our main results are shown in Figure 2 and 3 (please also see Figure 8, 9, and 10 in Appendix). In both these figures, we ranked the layers in the respective LLMs in descending order of their importance score. The left most point in the plots indicates that all layers of Mistral-7b and LLaMa2-7B are quantized in 4-bits; as we move to the right on x-axis, we quantize the next least important layer to 2 bits. For example, the overall bit size of 3.75 mentioned on the x-axis represents 28 (most important) layers in 4-bit and 4 (least important) layers in 2-bit. The horizontal red dotted lines indicate the performance of the entire model represented in 8bit precision; as shown, this performs very similarly to the full model in 4 bit precision (the top left most point in Figure 2 and 3). Because the gap between the full model in 8 bit vs. 4 bit precision was less than 1% across the majority of the datasets, we focus mostly on quantizing below 4-bit precision in our experiments.
4 to 2 bit 8 to 4 bit Models Layers GSM8K NQ_open GSM8K NQ_open 2 bits F1 F1 LIM Ordering LLama-Ins 5 7.5 15.0 10.5 36.6 10 1.5 6.7 13.5 37.9 15 0.5 3.1 11.0 35.4 Mistral-Ins 5 24.5 16.2 34.5 29.6 10 20.0 9.2 36.5 29.8 15 4.5 5.1 34.0 27.7 Z Ordering LLama-Ins 5 6.0 11.7 12.5 37.3 10 3.5 5.3 12.5 36.9 15 1.0 1.8 10.0 34.7 Mistral-Ins 5 25.0 15.4 37.0 28.5 10 10.5 8.7 34.5 30.1 15 1.0 2.2 34.0 29.3
Table 3: Performance comparison of LLaMa-instruct-7b and Mistral-instruct-7b across two generation tasks - GSM8K and Natural Questions open split.
We draw the following observations from our experiments:
  1. (1)
    Variable quantization is useful: The first key finding of our work is that a fixed quantization technique can be extended to a variable number of bits by quantizing different layers at different bits according to their importance. This allows LLMs to be fit in the exact memory requirement of the user while retaining more of the original performance. Overall, our method of layer-wise quantization, guided by layer importance, proves to be an efficient strategy for attaining adaptable precision bit levels.
  2. (2)
    Layer importance scoring is crucial: In the figures we compare layer ranking using our LIM and ZD importance scores with ranking using a reverse LIM and random ordering. As seen in fig. 2 and 3 the quantization of least important layers from 4-bit to 2-bit as per the LIM score ranking shows strong performance retention. In contrast, quantizing based on the reverse of LIM score shows much worse performance when most important layers are quantized to 2-bit, highlighting the strength of meaningful layer ranking. Further, LIM and ZD ranking performs substantially better than random ordering of layers baseline where we quantize layers to lower bits randomly. Lastly, LIM performs better than ZD, but the differences are not large. We consider this a success for ZD, which is simpler and does not require calibration data.
  3. (3)
    Layer-wise Quantization is useful until 3.0-3.25 bits: As shown in fig. 2 and 3, our first key observation is that quantization to 8 bit barely affects performance (red vs. purple line in each plot). While there is marginal drop in performance when models are quantized to 4 bits, we observed really noticeable drops only after the models are dropped below 3.0-3.25 bits on average using Quanto. For example, the bit size for which performance drops below 90% on Winogrande for Mistral-7b, LLaMa-7b, QWEN-7b, and LLaMa2-13b are 3.2, 3.1, 3.85, and 2.85 respectively.
    (a) Quantizing from 8 to 4 bits against pruning
    (b) Quantizing from 4 to 2 bits against pruning
    Figure 4: We compare quantization against pruning as a method to reduce the memory requirement of a model (LLaMa-2-7b here). One increment means two layers moved to lower quantization for the blue line (quantization), and one layer removed for the red and orange lines (pruning), thus reducing the same amount of memory. We show the average accuracy over MMLU, Winogrande, PIQA, and Hellaswag.
    Figure 5: Comparison of LLaMa2-7b quantized between 8 and 4 bits with LLaMa2-13b quantized between 4 and 2 bits to check when the performance intersects
  4. (4)
    Quantization is more useful for larger LLMs: Variable quantization of larger models (fig. 2) using our importance score shows much better retention of performance with lower quantization bit precision compared to moderately-sized LLMs such as LLaMa-7B (fig. 8 in appendix) and Mistral-7B (fig. 3). Further, as we go down to even smaller LLMs such as QWEN-1.8B (see fig. 10 in Appendix) with only 20 layers, we observed layer importance ranking to be not as effective. This observation aligns with many other previous works that have shown quantization to be substantially more effective for larger LLMs when compared to their smaller counterparts Jin et al. (2024).
  5. (5)
    Our method is applicable to different quantization techniques: Tables 1 and 2 summarize the overall results when quantizing individual layers with Quanto and GPT-Q, respectively. The tables show that our method can be coupled with any other quantization techniques. On average, GPT-Q leads to an average of 4% better accuracy than Quanto across all 5 tasks. Additionally, GPT-Q enables models such as LLaMa2-13B to be quantized down from 4 bits to 3.25 bits with less than a 3% loss in average accuracy, as seen in Table 2. Importantly, similar to Quanto and GPT-Q, our methodology of layer-wise quantization can be extended to any quantization technique.
  6. (6)
    Effect on generation tasks: — We also evaluate our approach of variable quantization of different layers on generation tasks. As shown in table 3, we observe substantial drop in performance on both GSM8K and NQ_open generation tasks when quantizing more layers in 2 bits. Importantly, the performance drop in these generation tasks is more drastic when compared to the average performance drop in classification tasks (table 1), emphasizing the need for more dedicated research in quantization for generation tasks.
6 Analyses
We present several analyses to further spotlight on benefits from variable layer-wise quantization.
6.1 Pruning vs. Quantization
We compare our variable quantization against variable pruning (using the same layer importance ranking) as an alternative for the same goal of reducing model memory requirement. In Figure 4(a), we show that for not extreme quantization levels, it is significantly better to move layers into lower quantization levels instead of removing them. For example, when 2 least important layers are removed resulting in remaining 30 layers (each in 8 bit) of LLaMa2-7b, the average performance drops to 62.7% (shown by red line denoting work by Gromov et al. (2024)). But on the quantization counterpart with same memory i.e., when 4 layers are quantized to 4 bit and remaining 28 layers are in 8 bit (shown by blue line), performance remains intact close to 66.8% as shown in fig. 4(a). When 12 layers are removed, the performance drops around 53% on average while having 8 layers in 8 bits and 24 layers quantized to 4 bits (shown by blue curve) to maintain the same size, average performance still remains intact and close to 66%. This highlights the important finding that quantization until 4 bits overall is a substantially more effective strategy compared to pruning for model compression.
On the other hand, in case of extreme levels of quantization (i.e., $<4-bits$) as shown fig. 4(b), it is better to plainly remove layers. For example, as shown in fig. 4(b) where the model is initially quantized with 4-bits, removing $>7$ layers results in better average performance when compared to quantizing $>14$ layers in 2 bits. Thus, when model compression is requires to be the equivalent of $<3bits$, pruning maybe the more effective strategy.
6.2 Quantizing Larger vs. Smaller LLMs
We further evaluate and compare feasibility of quantizing larger LLMs more drastically (i.e., $<4bits$ by quantizing less important layers in 2 bits and keeping more important in 4 bits) or quantizing smaller LLMs moderately (i.e., $<8bits$ by quantizing less important layers in 4 bits and keeping more important in 8 bits). As shown in fig. 5, we quantize LLaMa2-13B $<4bits$ and LLaMa2-7B $<8bits$ and compare them across different memory sizes. Our emperical finding suggests its more beneficial to quantize larger LLMs to smaller bits but only until a certain point, after which layer-wise quantizing smaller LLMs moderately ($6~{}bits$) shows better performance.
7 Conclusion
We introduced a simple, flexible quantization approach that quantizes different layers at different bits based on their importance. We presented two layer importance scoring techniques which when used to select more important layers for quantizing them in 4 bits and less important layers in 2 bits lead to strong performance retention across several LLMs even until 2.85 overall bit size. Our work presents several key practical findings such as layer-wise quantization is more effective for larger LLMs (with more number of layers), in the same memory setting; quantization is better than pruning until a certain bit precision level, text generation tasks are affected more with quantization, etc. Overall, our work introduces layer-wise quantization and presents detailed empirical findings for motivating future research in this direction.
8 Limitations
Limitations of our work are as follows:
  • Our experiments and results have focused more on quantization to lower bits i.e., $<4bits$ to highlight gains from layer-wise quantization based on their importance. As mentioned in section 5, we observed very similar performance between fp16 and 8-bit quantized model, minimal drop ($<$ 1% on average) in performance between 8-bit and 4-bit quantized model (please see table 4 in appendix). We did not observe meaningful changes in performance from layer-wise quantization between 8-bit and 4-bit because of such minute difference between their performances. Our study is limited to two level quantization i.e., more important layers in 4 bits and less important layers in 2 bits. One can also potentially apply three level of quantization i.e., most important layers in 8 bits, moderately important in 4 bits and least important in 2 bits. In our experiments, we observed three level quantization to be always worse than two level quantization for similar memory sizes (please see section A.1 in Appendix). .
  • We have proposed LIM and ZD scores for rating importance of each layer in LLM. These scores are very simple to implement, and only LIM needs a calibration corpus which is also an (readily available) unlabelled corpus. As future works, there can be more sophisticated measures to find importance of layers. For example, supervised methods that can show layer-wise impact on specific annotated NLP tasks Zhu et al. (2023a), in depth model interpretability approaches Singh et al. (2024); Sun et al. (2023), analyses of block of layers and their interactionsYang et al. (2024), etc. In this work, we limited and focused our study on LIM and ZD score to emphasize more on the main contribution of the paper - quantizing different layers at different bits as per layer importance. To the best of our knowledge, our work is the first to introduce variable quantization layer wise as per their importance and our study can be easily extended with more layer importance measuring scores in future work.
  • We presented comparison between pruning and quantization in section 6.1. A potential setup can also be combining pruning and quantization by removing few of the least important layers, quantizing remaining less important layers in lower 2 bits and keeping more important layers in 8 bits or above. Our work is primarily focused on introducing the idea of achieving variable quantization by quantizing different layers as per their importance. We leave exploration of combining quantization and pruning for future works.
9 Ethical Considerations
Our study is focused on post training quantization (PTQ) of LLMs. We have not finetuned LLMs for any specific task or data and have selected well established LLMs such as LLaMa, Mistral, and QWEN in our experiments. Our study also uses widely used evaluation datsets such as MMLU, HellaSwag, ARC easy, etc. that contain safe test cases. Thus, we believe our experiments do not contain any harmful cases and to the best of our knowledge, we did not observe any unsafe outputs in our evaluations.
Appendix A Appendix
Figure 6: We first rank the layers in an LLM (e.g., LLaMa2-13B) using an importance score (shown here is ranking based on our Layer Input Modification (LIM) score, see section 3.1).
A.1 Quantizing Layers Using 3 Levels
In all of our experiments, we have focused on two level quantization, i.e., either less important layers are quantized in 2 bits and more important layers in 4 bits or less important in 4 bits and more important in 8 bits. As a plausible variant, LLM layers can also be easily quantized using three levels, i.e., least important layers in 2 bits, moderately important in 4 bits, and the most important ones in 8 bits. In our study, we observed three level quantization almost always performs worse than two level quantization. We show three level quantization of LLaMa2-7b with fixed overall model bit size of 4bit. We first quantize all the 32 layers in 4 bits as shown by the leftmost bar of section A.1. We then convert two least important layers to 2 bits each and one most important layer to 8 bit, thus maintaining the overall bit size of the model to 4 bit. This is represented by second bar from the left in fig. 7. We repeat the same process of converting two more layers in 2bits and a more important layer to 8bit represented by the consecutive bars in fig. 7. As observed, three level quantization always performs worse than one level quantization when the target bit-level is the same, thus we don’t propose the technique as a way to achieve better performance for a set quantization level, but as a way to achieve a variable level of quantization while retaining maximum performance.
Figure 7: We compare different ways to achieve 4-bit quantization using three quantization levels. Each bar going from left to right represents adding one important layer in 8 bits and moving two less important layers to 2 bits, thus keeping an average of 4-bit quantization for all of the bars. Each bar having a value x on the x-axis represents the most important x layers in 8-bits, the least important 2*x in 2-bits and the rest in 4-bits.
Figure 8: Similarly to the other bit plots the graphs showcase the accuracy on four distinct data sets when quantizing LLaMa2-7b from full 4-bit quantization to 2-bit by moving less important layers in 2-bit quantization
Figure 9: Qwen2 quantized with quanto between 4 and 2 bits. All notations are same as in fig. 2.
Figure 10: Qwen1.5-1.8b quantized with quanto between 4 and 2 bits. All notations are same as in fig. 2.
Models Layers WNGD ARC PIQA HLSWG MMLU Average low-bits Accuracy BI Ordering LLaMa2-7B 5 69.06 75.88 77.69 57.05 41.28 64.2 10 68.82 76.3 77.42 57.11 41.96 64.3 15 68.82 76.3 77.42 57.11 41.96 64.3 Mistral-7B 5 73.87 81.01 80.9 61.13 58.56 71.1 10 73.95 80.3 80.95 61.28 58.39 71.0 15 73.79 80.42 80.79 61.17 58.14 70.9 LLaMa2-13B 5 72.29 79.33 79.16 60.15 50.49 68.3 10 71.74 79.08 79.32 59.96 50.53 68.1 15 71.74 79.37 79.32 59.96 50.55 68.2 Qwen-2-7B 5 70.71 79.2 80.03 58.66 68.06 71.3 10 70.95 79.2 79.97 58.44 67.68 71.2 15 68.82 78.32 79.81 58.26 67.66 70.6 Z Ordering LLaMa2-7B 5 68.82 76.13 77.91 57.02 40.85 64.1 10 68.66 75.92 77.63 57.09 40.6 64.0 15 68.27 75.54 77.42 57.09 39.7 63.6 Mistral-7B 5 73.87 80.76 80.9 61.24 58.62 71.1 10 74.19 80.47 81.12 61.03 58.2 71.0 15 74.42 80.47 80.84 60.89 58.31 71.0 LLaMa2-13B 5 72.29 79.54 78.94 60 50.54 68.3 10 72.21 79.58 79.16 59.99 50.51 68.3 15 72.05 79.46 79.37 59.98 50.59 68.3 Qwen-2-7B 5 71.58 79.2 79.76 59.2 69.44 71.8 10 71.5 79.08 80.35 58.82 69.01 71.8 15 71.58 78.61 79.92 58.54 68.65 71.5
Table 4: Accuracy results of different models across various tasks for 8bit and 4bit quantization.
A.2 LLaMa2 and QWEN plots
We show 4 bit to 2 bit variable layer-wise quantization for LLaMa2-7b and QWEN-7B in fig. 8 and fig. 9 respectively. All notations are same as in fig. 2. These curves were also generated on 2K evaluation instances from each of the datasets.
A.3 Commonality between layer importance
To the best of our knowledge, we are the first one to propose layer importance and utilize the importance order to quantize different layers at different bits. in Figure 6, we show the intensity bars below for each layer based on their importance for four different LLMs with different number of layers and sizes. As observed, there is a substantial pattern overlap of least important layers across multiple LLMs. We observe that first and the last layer are the most two important layers. Many of the important layers also tend to be the initial few set of layers. Lesser important layers (shown by block of layers with faded intensity) tend to be towards halfway of middle and end of the network. These observations suggest generalized patterns in layer importance across LLMs and pre-computed layer importance orders can be roughly utilized to quantize a wide variety of LLMs.
It is worth noting that (unpublished) concurrent works like Gromov et al. (2024) have presented an empirical finding that layers towards the end of the model can be removed except the last layer. The reverse order of layers indexes surprisingly has overlap with the importance order calculated with our LIM score but we believe our LIM score is more broadly applicable to any LLM where even the layers towards the end can be more important.
A.4 Results of 8bit to 4bit quantization on different datasets
We show results of quantizing models lower than 8 bit. Following our proposed methodology, the same techinque can be applied to have the more important layers in 8 bits and the least important ones in 4 bits. While this does still increase the performance that can be fit within a memory requirement, the results are not as major as the ones for the 4-2 bit range, thus we mainly focus on that range.